# Unmasking Trees for Tabular Data

### v4

@@ 3.4 Sales forecasting with uncertainty
4 Limitations
While UnmaskingTrees is overall state-of-the-art on the tabular imputation benchmark, MissForest still outperformed on the metrics based on Wasserstein distance to train and test dataset distributions. And Forest-Flow still won on vanilla generation benchmark (without any missingness). It remains to be seen whether a single method can be developed which wins on all scenarios and metrics. While BaltoBoTabPFN performed well on probabilistic prediction tasks, when used as a subroutine in UnmaskingTabPFN, it is very slow and experienced out-of-memory errors on the (Jolicoeur-Martineau et al., 2024b) benchmark on our machine. Further improvements either to it, or to how it is employed, are needed to make it practical for all but the smallest datasets.
@@ 5 Discussion and Related Work
Why is ForestDiffusion better at vanilla generative modeling, while UnmaskingTrees is better on missing data problems? We offer two speculative explanations. First, imputation is a conditional modeling scenario, except that you do not know the partition of the features into input features and output features a priori. One could address imputation by learning all possible $2^{D}$ conditional distributions, but this is impractical for large $D$, so one would prefer to learn a single joint distribution. Both autoregression and diffusion are ways of learning a joint distribution; because autoregression does so by learning conditional distributions, it is more suited to the conditional modeling imputation setting. Second, for missing data, diffusion has a train-inference gap: during training, observed features begin the reverse process from $\mathcal{N}(0,1)$; during inference for imputation, observed features begin the reverse process at their actual values. On the other hand, the advantages of diffusion modeling give it superiority when these problems can be avoided.
@@ 6 Conclusions
We proposed tree-based autoregressive modeling of tabular data, especially for data with missingness. For the subproblem of conditional probabilistic prediction of individual variables, we presented a hierarchical partitioning method with benefits over vanilla quantization and diffusion-based probabilistic prediction. We then considered each of these as meta-algorithms that enable pure in-context learning-based modeling using TabPFN as base classifier. We showed SotA results for imputation and for generation given data with missingness, and on probabilistic prediction for sales forecasting.
Appendix A Algorithmic descriptions of UnmaskingTrees and BaltoBot
The training algorithm for UnmaskingTrees is given in 1. The training and inference algorithms for BaltoBot are given in 2 and 3, respectively.
Algorithm 1 Unmasking Trees training
0:  dataset ${\mathbf{X}}\in{\mathbb{R}}^{N\times D}$; duplication factor $K$.
1:  {# Build self-supervised training set}
3:  for $k=1,\dots,K$ do
4:     for $n=1,\dots,N$ do
5:        Draw random permutation $\sigma$ from ${\mathcal{U}}(G_{D})$
6:        Set $\bm{x}:={\mathbf{X}}_{n,:}$ and $\bm{y}:={\mathbf{X}}_{n,:}$.
7:        for $d=1,\dots,D$ do
[…]
10:        end for
11:     end for
12:  end for
13:  {# Train conditional generation models}
14:  for $d=1,\dots,D$ do
15:     if feature $d$ in ${\mathbf{X}}$ is a continuous feature then
16:        Run BaltoBot with $([{\mathbf{X}}_{\textrm{train}}]_{:,j\neq d},[{\mathbf{Y}}_{\textrm{train}}]_{:,d})$.
17:     else
18:        Train XGBClassifier on $([{\mathbf{X}}_{\textrm{train}}]_{:,j\neq d},[{\mathbf{Y}}_{\textrm{train}}]_{:,d})$.
19:     end if
20:  end for
Algorithm 2 BaltoBot training
1:  if H = 0 or unique$({\mathbf{y}})=C$ for some constant $C$ then
2:     Save $\texttt{bounds}:=(\min({\mathbf{y}}),\max({\mathbf{y}}))$.
3:  else
4:     Obtain split point $p$ from KDI quantization on ${\mathbf{y}}$.
8:  end if
Algorithm 3 BaltoBot inference
0:  input query ${\mathbf{x}}\in{\mathbb{R}}^{D}$; trained BaltoBot model.
1:  if bounds is defined then
2:     Sample uniformly from $U(\texttt{bounds})$.
3:     Return.
4:  else
5:     Obtain prediction from XGBoost binary classifier.
6:     if prediction = left-child then
7:        Run inference on “left-child” BaltoBot with input query ${\mathbf{x}}$.
8:     else if prediction = right-child then
10:     end if
11:  end if
Appendix B Ablation experiment with imputation - raw scores
@@
Appendix C Full dataset-level results
Full imputation results are in Table 7. Full generation results are in Table 9. Timing results are in Table 10, and depicted in Figure 5. Our method is relatively efficient at both imputation and generation. The datasets on which we are slowest for imputation are Libras (1976 seconds) and Bean (1929 seconds), on our ancient 2015 iMac with 16Gb RAM. On Libras, ForestVP imputation took 12439 seconds (without RePaint) and 14715 seconds (with RePaint); on Bean, ForestVP took 898 seconds (without RePaint) and 1318 seconds (with RePaint), on their cluster of 10-20 CPUs with 64-256Gb of RAM. The datasets on which we are slowest for generation are also Libras (2987 seconds) and Bean (4346 seconds). On Libras, ForestFlow generation took 9481 seconds and ForestVP took 9042 seconds; on Bean, ForestFlow took 869 seconds and ForestVP took 947 seconds, once again on their much more powerful computing cluster.
@@ Appendix C Full dataset-level results
Dataset MinMAE $\downarrow$ AvgMAE $\downarrow$ $P_{bias}\downarrow$ $Cov_{rate}\uparrow$ $W_{train}\downarrow$ $W_{test}\downarrow$ Variance $\uparrow$ MAD (mean) $\uparrow$ MAD (median) $\uparrow$ $R^{2}$uparrow F1 $\uparrow$
@@
breast cancer diagnostic 3.96e-02 5.16e-02 0.00e+00 0.00e+00 3.10e-01 1.85e+00 1.21e-03 1.01e-01 8.73e-02 0.00e+00 9.59e-01
connectionist bench vowel 5.30e-02 9.39e-02 0.00e+00 0.00e+00 1.88e-01 7.25e-01 5.53e-03 2.48e-01 2.14e-01 0.00e+00 6.64e-01

### v5

@@ 3.4 Sales forecasting with uncertainty
4 Limitations and Future Work
Limitations
While UnmaskingTrees leads overall on the tabular imputation benchmark, MissForest still outperformed on the metrics based on Wasserstein distance to train and test dataset distributions. And Forest-Flow still won on vanilla (i.e. no missingness) generation benchmark. It remains to be seen whether a single method can be developed which wins on all scenarios and metrics.
Our proposed BaltoBot method would benefit from a more principled method for selection of the meta-tree height $H$. Unlike with standard flat quantization where having a large number of bins can cause one to make catastrophically wrong predictions, BaltoBot “knows” proximities among meta-tree leaves. This means that making $H$ bigger tends not to cause major errors, so it’s better to err on the side of larger $H$ rather than small $H$. Still, the main drawbacks with increasing $H$ are that (1) training takes longer, and (2) imputations and generations are less diverse. Deeper theoretical analysis of these trade-offs and a more principled approach for choosing the height would improve the ease-of-use and potentially the performance of our method.
While BaltoBoTabPFN performed well on probabilistic prediction tasks, when used as a subroutine in UnmaskingTabPFN, it is very slow and experienced out-of-memory errors on the (Jolicoeur-Martineau et al., 2024b) benchmark on our machine. Further improvements either to it, or to how it is employed, are needed to make it practical for all but the smallest datasets.
With regards to scalability, the runtime complexity of UnmaskingTrees training is cubic in terms of feature dimensionality $D$. This compares to quadratic complexity for ForestDiffusion; this would also be the complexity for single-order autoregressive modeling, which would support only generation rather than arbitrary imputation. This means that our approach is primarily applicable for small-sized tabular datasets, or for scenarios where inference time is more important than training time.
Related to the above remarks, we would like to emphasize that our proposed approach is aimed at and evaluated on smaller-sized tabular datasets. It is also evaluated via “out-of-the-box” performance, being aimed at users lacking the resources for large deep learning models or hyperparameter optimization. For users with access to larger tabular datasets and more extensive computing resources, recent deep learning methods like Tabsyn (Zhang et al., 2024) would be expected to perform better.
Future Work
Reducing the training complexity with respect to the number of features is a key next step. One possibility would be to use an optimized, rather than random, selection of feature orderings at training time (Shih et al., 2022). Another possibility would be to use multi-output trees to train a single XGBoost model for all BaltoBot tree nodes and all features, similarly to a recently-proposed approach for speeding up ForestDiffusion (Cresswell and Kim, 2024). In addition to improving scalability, BaltoBot’s core idea of binary partitioning could be combined with deep learning approaches. For example, one might equip a neural network with a hierarchical softmax head (Morin and Bengio, 2005) for modeling continuous outputs without losing proximity information among bins. Finally, future work could theoretically analyze BaltoBot, empirically evaluate it on uncertainty quantification and probabilistic forecasting tasks, and extend it to multiple dimensions.
@@ 5 Discussion and Related Work
Why is ForestDiffusion better at vanilla generative modeling, while UnmaskingTrees is better on missing data problems? We offer two speculative explanations. First, imputation is a conditional modeling scenario, except that you do not know the partition of the features into input features and output features a priori. One could address imputation by learning all possible $2^{D}$ conditional distributions, but this is impractical for large $D$, so one would prefer to learn a single joint distribution. Both autoregression and diffusion are ways of learning a joint distribution; because autoregression does so by learning conditional distributions, it is more suited to the conditional modeling imputation setting. Second, for missing data, diffusion has a train-inference gap: during training, observed features begin the reverse process from $\mathcal{N}(0,1)$; during inference for imputation, observed features begin the reverse process at their actual values. On the other hand, the advantages of diffusion modeling (no quantization error, holistic generation, needing only an estimated score function rather than well-calibrated conditional distributions) give it superiority when these problems can be avoided.
@@ 6 Conclusions
We proposed tree-based autoregressive modeling of tabular data, especially for data with missingness. For the subproblem of conditional probabilistic prediction of individual variables, we presented a hierarchical partitioning method with benefits over vanilla quantization and diffusion-based probabilistic prediction. We then considered each of these as meta-algorithms that enable pure in-context learning-based modeling using TabPFN as base classifier. On a benchmark for out-of-the box performance on tabular data, we showed leading results for imputation and state-of-the-art results for generation given data with missingness.
Appendix A Ablation experiment with imputation - raw scores
@@
Appendix B Full dataset-level results
Full imputation results are in Table 7. Full generation results are in Table 8. Timing results are in Table 10, and depicted in Figure 5. Our method is relatively efficient at both imputation and generation. The datasets on which we are slowest for imputation are Libras (1976 seconds, $N=360$, $D=90$) and Bean (1929 seconds, $N=13611$, $D=16$), on our ancient 2015 iMac with 16Gb RAM. On Libras, ForestVP imputation took 12439 seconds (without RePaint) and 14715 seconds (with RePaint); on Bean, ForestVP took 898 seconds (without RePaint) and 1318 seconds (with RePaint), on their cluster of 10-20 CPUs with 64-256Gb of RAM. The datasets on which we are slowest for generation are also Libras (2987 seconds) and Bean (4346 seconds). On Libras, ForestFlow generation took 9481 seconds and ForestVP took 9042 seconds; on Bean, ForestFlow took 869 seconds and ForestVP took 947 seconds, once again on their much more powerful computing cluster.
@@ Appendix B Full dataset-level results
Dataset MinMAE $\downarrow$ AvgMAE $\downarrow$ $P_{bias}\downarrow$ $Cov_{rate}\uparrow$ $W_{train}\downarrow$ $W_{test}\downarrow$ Var $\uparrow$ MAD (mean) $\uparrow$ MAD (med) $\uparrow$ $R^{2}$uparrow F1 $\uparrow$
@@
breast cancer 3.96e-02 5.16e-02 0.00e+00 0.00e+00 3.10e-01 1.85e+00 1.21e-03 1.01e-01 8.73e-02 0.00e+00 9.59e-01
connectionist vowel 5.30e-02 9.39e-02 0.00e+00 0.00e+00 1.88e-01 7.25e-01 5.53e-03 2.48e-01 2.14e-01 0.00e+00 6.64e-01